# In-Context Former: Lightning-fast Compressing Context for Large Language Model

### v1

@@ Abstract
Xiangfeng Wang1, Zaiyi Chen1, Zheyong Xie1, Tong Xu1thanks: Corresponding author., Yongyi He1, Enhong Chen1 1University of Science and Technology of China {xf9462, czy6516, xiezheyong, vagabond}@mail.ustc.edu.cn, {tongxu, cheneh}@ustc.edu.cn
@@ 1 Introduction
Figure 1: Compressing long contexts into short soft prompts to improve inference efficiency.
To better avoid modifications to the LLM structure, a more intuitive approach is to introduce a preliminary context compression process. These methods are based on a core assumption: most natural language texts contain redundant information, which makes context compression feasible. In early exploration, Mu et al. (2024) have attempted to compress the instructions into short soft prompts. This method offers a novel perspective but still has limitations in long context compression. Later works Chevalier et al. (2023); Ge et al. (2024) aim to further extend compression abilities for document-level long contexts, and achieved considerable results. As illustrated in Figure 1, these methods design compression models to condense lengthy contexts into short, context-rich soft prompts, which then serve as substitutes for the original context when input into the LLM. However, these methods still suffer the issue of expensive time costs during the compression process. This limitation restricts their application in real-time compression scenarios, such as compressing retrieved  Guu et al. (2020) or real-time Internet documents  Asai et al. (2023) immediately.
@@
Driven by this motivation, we design an efficient context compression model, the In-Context Former (IC-Former), which aims at optimizing resource consumption during the compression of long context in existing models. This model is based on two assumptions regarding semantic content compression: (1) Word embeddings already contain sufficient semantic information Mikolov et al. (2013); Tache et al. (2021), suggesting that additional interactions may not be necessary prior to the extraction process. (2) Learnable tokens within an elaborate structure can effectively aggregate information to a certain extent Chevalier et al. (2023); Ge et al. (2024). Based on these assumptions, we creatively discard the costly self-attention interaction of text content in previous models. Instead, we leverage the efficiency of the cross-attention mechanism for information extraction. This innovative strategy ensures that the computational overhead of compression grows linearly with the context length within the compression range, significantly enhancing compression efficiency compared to the previous methods.
@@ 3.2 In-Context Former
Then we employ the cross-attention mechanism to condense contextual information, as this approach has been empirically validated effective. Li et al. (2023); Ye et al. (2023); Zhu et al. (2023).
@@
In the Eq.3.2, $\theta=\theta_{base}^{-\frac{2}{h}}$ where $\theta_{base}$ is a hyper-parameter and $h$ is the hidden size and assumed to be even. We restate Eq.1 & 2 as follows:
@@
We allocate positional embeddings as if placing the digest tokens subsequent to the context tokens as demonstrated in Eq.7 & 8.
@@ 4.2.1 Compression & Inference Efficiency
Table 4: Evaluation results on PwC test set. The first row of the table compares the performance of our method with other baseline models, and the performance ratio means the ratio of our IC-Former to the ICAE. The second row demonstrates the performance variations when different compression strategies are implemented, where "64+64" represents a divide-and-conquer approach. The third row reveals the impact of ablation pre-training on performance.
@@ 4.2.3 Performance on Downstream Task
Figure 6: Local attention map in the last layer of IC-Former. The horizontal axis represents context tokens acting as key and the vertical axis represents digest tokens acting as query. For complete attention map, see Appendix E.
@@ 4.3 Analysis
A large language model (LLM) is a computational model notable for its ability to achieve general-purpose language generation and other natural language processing tasks such as classification. Based on language models, LLMs acquire these abilities by learning statistical relationships from vast amounts of text during a computationally intensive self-supervised and semi-supervised training process. LLMs can be used for text generation, a form of generative AI, by taking an input text and repeatedly predicting the next token or word. LLMs are artificial neural networks that utilize the transformer architecture, invented in 2017. The largest and most capable LLMs, as of June 2024, are built with a decoder-only transformer-based architecture, which enables efficient processing and generation of large-scale text data. Larger models such as GPT-3 have demonstrated the ability to achieve similar results through prompt engineering, which involves crafting specific input prompts to guide the model’s responses.
Table 6: The context tokens that are most attended to by digest tokens across layers. The color of each token is determined by the layer when it is initially attended to. Red, cyan, and blue denote the first, second and third layer respectively. Gray indicates tokens that are never attended to.
@@
As illustrated in Table 6, it can be observed that in the first layer, digest tokens mainly focus on prepositions, articles, be-verb, and punctuation marks. As we proceed to the second layer, digest tokens start to extend their focus to verbs, nouns, adjectives, and adverbs. The third layer continues this trend based on the second layer, further broadening the range of grammatical categories of tokens covered, encompassing a more extensive context. This implies that IC-Former might rely on semantic structures to understand and compress context effectively.

### v2

@@ Abstract
Xiangfeng Wang, Zaiyi Chen, Tong Xuthanks: Corresponding author., Zheyong Xie, Yongyi He, Enhong Chen University of Science and Technology of China {xf9462, czy6516, xiezheyong, vagabond}@mail.ustc.edu.cn, {tongxu, cheneh}@ustc.edu.cn https://github.com/wonderful9462/IC-Former
@@ 1 Introduction
Figure 1: Compressing long contexts into short soft prompts (vectors in embedding space) to improve inference efficiency.
To better avoid modifications to the LLM structure, a more intuitive approach is to introduce a preliminary context compression process. These methods are based on a core assumption: most natural language texts contain redundant information, which makes context compression feasible. In early exploration, Mu et al. (2024) have attempted to compress the instructions into short soft prompts. This method offers a novel perspective but still has limitations in long context compression. Later works Chevalier et al. (2023); Ge et al. (2024) aim to further extend compression abilities for document-level long contexts, and achieved considerable results. As illustrated in Figure 1, these methods design compression models to condense lengthy contexts into short, context-rich soft prompts, which then serve as substitutes for the original context when input into the LLM. However, these methods still suffer the issue of expensive time costs during the compression process. This limitation restricts their application in real-time compression scenarios, such as compressing retrieved  Guu et al. (2020); Lyu et al. (2024) or real-time Internet documents  Asai et al. (2023) immediately.
@@
Driven by this motivation, we design an efficient context compression model, the In-Context Former (IC-Former), which aims at optimizing resource consumption during the compression of long context in existing models. This model is based on two assumptions regarding semantic content compression: (1) Word embeddings already contain sufficient semantic information Mikolov et al. (2013); Tache et al. (2021), suggesting that interactions between embeddings may not be necessary prior to the extraction process. (2) Learnable tokens within an elaborate structure can effectively aggregate information to a certain extent Chevalier et al. (2023); Ge et al. (2024). Based on these assumptions, we try to discard the costly self-attention interaction of text content in previous models. Instead, we leverage the efficiency of the cross-attention mechanism for information extraction. This innovative strategy ensures that the computational overhead of compression grows linearly with the context length within the compression range, significantly enhancing compression efficiency compared to the previous methods.
@@ 3.2 In-Context Former
Then we employ the cross-attention mechanism to condense contextual information, as this approach has been empirically validated effective in multimodal information extraction. Li et al. (2023); Ye et al. (2023); Zhu et al. (2023); Bai et al. (2023).
@@
In the Equation.3.2, $\theta=\theta_{base}^{-\frac{2}{h}}$ where $\theta_{base}$ is a hyper-parameter and $h$ is the hidden size and assumed to be even. We restate Equation.1 & 2 as follows:
@@
We allocate positional embeddings as if placing the digest tokens subsequent to the context tokens as demonstrated in Equation.7 & 8.
@@ 4.2.1 Compression & Inference Efficiency
Table 4: Evaluation results on PwC test set. The first four rows of the table compare the performance of our method with other baseline models, and the performance ratio means the ratio of our IC-Former to the ICAE. The second three rows demonstrate the performance variations when different compression strategies are implemented, where "64+64" represents a divide-and-conquer approach. The last row reveals the impact of ablation pre-training on performance.
@@ 4.2.3 Performance on Downstream Task
Figure 6: A part of attention map in the last layer of IC-Former. The horizontal axis represents context tokens acting as key and the vertical axis represents digest tokens acting as query. For complete attention map, see Appendix E.
@@ 4.3 Analysis
A large language model (LLM) is a computational model notable for its ability to achieve general-purpose language generation and other natural language processing tasks such as classification. Based on language models, LLMs acquire these abilities by learning statistical relationships from vast amounts of text during a computationally intensive self-supervised and semi-supervised training process. LLMs can be used for text generation, a form of generative AI, by taking an input text and repeatedly pGreenicting the next token or word. LLMs are artificial neural networks that utilize the transformer architecture, invented in 2017. The largest and most capable LLMs, as of June 2024, are built with a decoder-only transformer-based architecture, which enables efficient processing and generation of large-scale text data. Larger models such as GPT-3 have demonstrated the ability to achieve similar results through prompt engineering, which involves crafting specific input prompts to guide the model’s responses.
Table 6: The context tokens that are most attended to by digest tokens across layers. The color of each token is determined by the layer when it is initially attended to. Green, blue, and red denote the first, second and third layer respectively. Gray indicates tokens that are never attended to.
@@
As illustrated in Table 6, it can be observed that in the first layer, digest tokens mainly focus on prepositions, articles, be-verb, and punctuation marks. As we proceed to the second layer, digest tokens start to extend their focus to verbs, nouns, adjectives, and adverbs. The third layer continues this trend based on the second layer, further broadening the range of grammatical categories of tokens covered, encompassing a more extensive context. This implies that IC-Former might rely on semantic structures to compress context effectively.
@@ 5 Conclusion
Acknowledgments
This work was supported by the grants from National Natural Science Foundation of China (No.62222213, 62072423).